Course Title: ZERO-KNOWLEDGE Semesteromfang/timetal: 0.2 semester / 2 hours of lectures per week, plus 2 office

hours each week.

Prerequisites: A course in cryptology. Aims and Objectives: The purpose of this course is to give the paricipants an un-

derstanding of interactive proof systems and zero-knowledge, so that they can design protocols which are secure. Some complexity-theoretic results which have come about through research on interactive proofs may also be discussed.

Course Description: Interactive proof systems, zero-knowledge proofs of membership and proofs of knowledge, computationally convincing arguments, multiple-prover interactive proofs.

Grading method: Undervisningsform: Reading course Schedule: To be announced. Lecturer: Joan Boyar. Literature: Articles. Remarks: This course is also approved as an elcective course for the

master’s degree.